OpenStudy (anonymous):

Choose the correct simplification of the expression 4 over y to the negative 3

OpenStudy (rosedewittbukater):

Are there choices?

OpenStudy (anonymous):

4y3 4 over y to the 3rd power y to the 3rd power over 4 This expression is already simplified.

OpenStudy (rosedewittbukater):

is it \[\frac{ 4 }{ y^{-3} }\] Or \[\left( \frac{ 4 }{ y } \right)^{-3}\]

OpenStudy (rosedewittbukater):

I'll assume it's the first expression I put. There can't be any negative exponents in the denominator (the part under the fraction line). Usually if there is a number with a negative exponent, you would get rid of the negative by making the number a fraction. For example, \[5 ^{-3}\]would become \[\frac{ 1 }{ 5^ {3} }\]Doing that would get rid of the negative. Since the negative is in the denominator, you do the opposite to get rid of the negative. Do you know what to do?

OpenStudy (anonymous):

Yes, thank you. It was very helpful.

OpenStudy (rosedewittbukater):

Glad to help